calculate the probabilty of randomly guessing the given number of correct answers on a 30-question multiple-choice exam that has choices A, B,C,and D for each question.
#19) 2

Answer by edjones(8007)

The chance of guessing exactly 2 answers correctly out of 30 questions.
Let b=1/4, the chance of guessing a question correctly and a=3/4, guessing it wrong.
The coefficient of a^28b^2 is nCr=30C2=435
435*(3/4)^28*(1/4)^2=.008631...
.
Ed